# Inverse analysis of granular flows using differentiable graph neural network simulator

### v2

@@ 7 Data and code availability
The GNS code (Kumar and Vantassel, 2022) is available under the MIT license on GitHub (https://github.com/geoelements/gns). The training dataset and trained models are published under CC-By license on DesignSafe Data Depot (Choi and Kumar, 2023b).
@@ 8 Acknowledgment
This material is based upon work supported by the National Science Foundation under Grant No.#2103937. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Performance of GNS
As a benchmark test for Flow2D, we show the performance of GNS using the granular column collapse. Granular column collapse (Lajeunesse et al., 2005; Utili et al., 2015; Lube et al., 2005; Kumar, 2015) is an experiment to study the dynamics of granular flows in a controlled setting. A granular column of initial height $H_{0}$ and length $L_{0}$ is placed on a flat surface, and it is allowed to collapse under gravity. The dynamics of granular column collapse is majorly governed by the initial aspect ratio of the column ($a=H_{0}/L_{0}$). The previous study Choi and Kumar (2023a) proves that GNS can accurately predict different flow dynamics of granular columns with various aspect ratios beyond which the data the model is trained. In this paper, we briefly show the prediction performance of the GNS used in this study on a short column ($a=0.5$) with different friction angles ($\phi=21\degree$ and $42\degree$), whose aspect ratio and friction angles are not seen during the training.
Figure 1 shows the evolution of granular flow for the short column ($a=0.5$) with normalized time ($t/\tau_{c}$) simulated by GNS and MPM. Here, we assume the MPM result is the ground truth. $t$ is physical time, and $\tau_{c}$ is the critical time defined as the time required for the flow to fully mobilize. $\tau_{c}$ is defined as $\sqrt{H_{0}/g}$, where $g$ is the gravitational acceleration. Figure 0(a) and fig. 0(b) is the result with $\phi=21\degree$ and $\phi=42\degree$. Each row of the figure corresponds to flow (1) at the time before initiation of collapse, (2) $t/\tau_{c}=1.0$, (3) $2.5$, and (4) at the last time step ($k=380$) when the flow reaches static equilibrium and forms the final deposit.
@@
Figure 1: Evolution of material point flow with normalized time for GNS and MPM for the short column with a = 0.5: (a) $\phi=21\degree$, (b) $\phi=42\degree$. The color represents the magnitude of the displacement. Each row corresponds to a column before the initiation of flow, $t/\tau_{c}=1.0$, $t/\tau_{c}=2.5$, and the final deposit at the last timestep.
@@
Figure 2 shows an example of GNS prediction trained on Obstacle3D datasets. The test configuration is outside the training data distribution as summarized in 1. Specifically, the simulation domain is four times larger with 2.2 times more material points compared to the training data. The maximum length of the initial granular mass in our test is 1.7 m, while that of training data is 0.7 m. Additionally, we test the GNS with five baffles, versus one to three baffles in the training data.
@@
From the initial state to t=0.125 s (fig. 2a-b), the granular debris propagates downstream uniformly and impacts the first baffle row. Upon hitting this row, the baffles obstruct the sides of the flow while the material between the baffles proceeds towards the next row. Concurrently, material deposits and builds up upstream of the baffles, reaching almost the height of the baffle. From t=0.125 s to t=0.325 s (fig. 2b-c), the flow impacts the next baffle row and diverges into four granular jets through the open area with roughly symmetric shapes. Less material deposition forms upstream of the last baffle row compared to the first row. Beyond t=0.325 s (fig. 2c-d), the spreading of the grains decelerates due to basal friction and finally reaches static equilibrium around t=0.875 s. At this stage, the runout error between GNS and MPM is 0.87%. The GNS rollout successfully replicates the overall kinematics including complex baffle interactions.
@@
Figure 2: Evolution of flow interacting with baffles for GNS and MPM from initial condition to the final deposit at the last timestep. The simulation domain is $1.8\times 0.8\times 1.8\ m$ and the initial geometry of the granular mass is $0.35\times 0.25\times 1.4\ m$. The barrier size is $0.15\times 0.30\times 0.15\ m$. The center location of the three baffles in the first row are at $(x=0.76,\ z=0.36)\ m$, $(x=0.76,\ z=1.00)\ m$, and $(x=0.76,\ z=1.64)\ m$, and the two baffles in the second row are at $(x=1.26,\ z=0.68)\ m$ and $(x=1.26,\ z=1.32)\ m$

### v3

@@ 7 Data and code availability
The code for GNS (Kumar and Vantassel, 2022) and inverse problems are available under the MIT license on GitHub (https://github.com/geoelements/gns and https://github.com/geoelements/gns-inverse-examples). The training dataset, trained models, and the data used for the inverse analysis are published under CC-By license on DesignSafe Data Depot (Choi and Kumar, 2023b, 2024).
@@ 8 Acknowledgment
This material is based upon work supported by the National Science Foundation under Grant No.#2103937 and #2229702. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Performance of GNS
As a benchmark test for Flow2D, we show the performance of GNS using the granular column collapse. Granular column collapse (Lajeunesse et al., 2005; Utili et al., 2015; Lube et al., 2005; Kumar, 2015) is an experiment to study the dynamics of granular flows in a controlled setting. A granular column of initial height $H_{0}$ and length $L_{0}$ is placed on a flat surface and allowed to collapse under gravity. The dynamics of granular column collapse is majorly governed by the initial aspect ratio of the column ($a=H_{0}/L_{0}$). The previous study Choi and Kumar (2023a) proves that GNS can accurately predict different flow dynamics of granular columns with various aspect ratios beyond which the data the model is trained. In this paper, we briefly show the prediction performance of the GNS used in this study on a short column ($a=0.5$) with different friction angles ($\phi=21\degree$ and $42\degree$), whose aspect ratio and friction angles are not seen during the training.
Figure 1 shows the evolution of granular flow for the short column ($a=0.5$) with normalized time ($t/\tau_{c}$) simulated by GNS and MPM. Here, we assume the MPM result is the ground truth. $t$ is physical time, and $\tau_{c}$ is the critical time, defined as the time required for the flow to mobilize fully. $\tau_{c}$ is defined as $\sqrt{H_{0}/g}$, where $g$ is the gravitational acceleration. Figure 1(a) and fig. 1(b) is the result with $\phi=21\degree$ and $\phi=42\degree$. Each row of the figure corresponds to flow (1) at the time before initiation of collapse, (2) $t/\tau_{c}=1.0$, (3) $2.5$, and (4) at the last timestep ($k=380$) when the flow reaches static equilibrium and forms the final deposit.
@@
Figure 1: Evolution of material point flow with normalized time for GNS and MPM for the short column with a = 0.5: (a) $\phi=21\degree$, (b) $\phi=42\degree$. The color represents the magnitude of the displacement. Each row corresponds to a column before the flow initiation, $t/\tau_{c}=1.0$, $t/\tau_{c}=2.5$, and the final deposit at the last timestep.
@@
Figure 2 shows an example of GNS prediction trained on Obstacle3D datasets. The test configuration is outside the training data distribution as summarized in 1. Specifically, the simulation domain is four times larger, with 2.2 times more material points than the training data. The maximum length of the initial granular mass in our test is 1.7 m, while that of training data is 0.7 m. Additionally, we test the GNS with five baffles, versus one to three baffles in the training data.
@@
From the initial state to t=0.125 s (fig. 2a-b), the granular debris propagates downstream uniformly and impacts the first baffle row. Upon hitting this row, the baffles obstruct the sides of the flow while the material between the baffles proceeds towards the next row. Concurrently, material deposits and builds up upstream of the baffles, reaching almost the height of the baffle. From t=0.125 s to t=0.325 s (fig. 2b-c), the flow impacts the next baffle row and diverges into four granular jets through the open area with roughly symmetric shapes. Less material deposition forms upstream of the last baffle row than the first. Beyond t=0.325 s (fig. 2c-d), the spreading of the grains decelerates due to basal friction and finally reaches static equilibrium around t=0.875 s. At this stage, the runout error between GNS and MPM is 0.87%. The GNS rollout successfully replicates the overall kinematics, including complex baffle interactions.
@@
Figure 2: Evolution of flow interacting with baffles for GNS and MPM from initial condition to the final deposit. The simulation domain is $1.8\times 0.8\times 1.8\ m$, and the initial geometry of the granular mass is $0.35\times 0.25\times 1.4\ m$. The barrier size is $0.15\times 0.30\times 0.15\ m$. The center location of the three baffles in the first row is at $(x=0.76,\ z=0.36)\ m$, $(x=0.76,\ z=1.00)\ m$, and $(x=0.76,\ z=1.64)\ m$, and the two baffles in the second row are at $(x=1.26,\ z=0.68)\ m$ and $(x=1.26,\ z=1.32)\ m$.
To quantitatively compare the results from the GNS and MPM, the evolution of the runout distance is measured (fig. 3a). Here, runout is defined as the flow’s maximum travel distance compared to the initial granular mass geometry. As flow initiates, the runout rapidly increases until 0.325. Subsequently, the runout gradually decelerates and eventually stops due to the interaction with the baffles and frictional dissipation. GNS accurately predicts the runout evolution simulated by MPM despite the simulation including more barriers and a larger domain than the training data.
The upstream depth evolution of the granular mass deposition behind the baffles is also measured (fig. 3a). The upstream flow depth refers to the depth measured upstream of the baffle array when the flow impacts the baffles. As the runout proceeds, the flow faces the first and then a second row of baffles. Upon impact with the first baffles, the upstream depth spikes close to the baffle height (0.3 m) due to the deposition of the granular mass. It then decreases slightly over time as the flow around the baffles reaches equilibrium. When the flow hits the next baffle row, the upstream depth surges again and gradually decreases. GNS accurately replicates the overall upstream depth change trend with minor errors.
Figure 3b shows the energy evolution. The initial granular flow shows high kinetic and potential energy. Subsequent granular flow and its blockage due to the baffles cause the kinetic and potential energy drop. GNS precisely captures these energy evolution trends. The results from fig. 3 suggest that GNS successfully learns the granular flow dynamics with obstacle interactions.
((a))
((b))
Figure 3: Quantitative comparison between MPM and GNS: (a) Runout and upstream depth evolution with time, (b) potential, kinetic and dissipation energy ($E_{p}$, $E_{k}$, and $E_{d}$) evolution with time normalized by $E_{0}$, where $E_{0}$ is total energy at $t=0$.
[…]
Figure 4: Normalized runout error between MPM and GNS evaluated at different friction angles $\phi$. The final deposit at points (a), (b), and (c) is shown below in the plot for the normalized runout error.
Appendix B Training data
Figure 1: An example of training data for 2D GNS (Flow2D).
Figure 2: An example of training data for 3D GNS (Obstacle3D).